OpenStudy (anonymous):

7y=y-42 solve step by step with work

OpenStudy (anonymous):

7y=y-42 Subtract y from both sides 6y=-42 Divide both sides by 6 y=-7

OpenStudy (merchandize):

7y-y=-42 6y=-42 y=-42/6 y=-7